OpenStudy (anonymous):

Simplify: 6sqrt2-3sqrt2 . Use the Equation Editor to write your answer in simplest radical form

OpenStudy (unklerhaukus):

\[6\sqrt 2-3\sqrt2=(6-3)\sqrt2=\cdots\sqrt2\]

OpenStudy (unklerhaukus):

\[=3\sqrt 2\]